OpenStudy (zephyr141):

according to the position vs. time graph below, for what duration is velocity positive?

OpenStudy (zephyr141):

 A. over the whole duration B. from t=0 until t=4 C. from t=2 until t=14 D. between t=4 and t=8

OpenStudy (zephyr141):

so i'm going to go with and C. i mean since velocity has a direction so it would be negative if it's going in the negative x direction right?

OpenStudy (danjs):

Velocity is the slop of the line on the position vs time graph. A positive velocity and positive slope occurs at time t=0 to time t=4. If the line has zero slope, (the horizontal portion), the object is at rest. ie. time is passing while the object stays at the same x position.

OpenStudy (zephyr141):

hmmm... cripes. i need to study more. thanks.

OpenStudy (danjs):

velocity is a change in position v= delta x / delta t [rise over run on the x vs t graph]

OpenStudy (zephyr141):

ok. I see now.

OpenStudy (zephyr141):

4s to 8s it's not moving. 8s to 20s it has negative velocity

OpenStudy (danjs):

correct. And to calculate the velocity, you have to calculate the slope of the line segments. (horizontal part = 0)

OpenStudy (danjs):

From 0 to 4 seconds v = delta x / delta t = [X2-X1]/[t2-t1] v= [6-(-6)]/[4-0] = 12/4 v=(+) 3 m/s

OpenStudy (zephyr141):

yeah i just put that into my calculator too. got the same as you :)

OpenStudy (danjs):

cool. one step further... the acceleration is zero for this graph if you plot your velocity vs time from this graph you get constant velocities over each interval. so the change in the velocities is zero

OpenStudy (zephyr141):

so because it's linear on all portions of the graph it's constant velocity? if it was parabolic it would have different acceleration?